solid NH4I on rapid heating in a closed vessel at357 degree celcius develops a const.pessue of 275 mm hg owing to partial decompositn of NH4I into NH3 and HI but pressure gradually increases further [when excess solid residue remains in vessel] owing to decompositn of HI.calculate final pressure developed at equlibrium.
NH4I(s)forms NH3(g) + HI(g)
2HI(g) forms H2(g)+I2(g)  Kc=0.065 at 357 degree celcius.

Grade:12

2 Answers

Ashutosh Sharma
8 years ago
NH4I(g) NH3(g) + HI(g) ------(1)
using Kp and Kc relation Kp can be found out or for HI division Kp is 0.015.
for this eqn 2p=275 –>p=137.5
So kpfor eq. (1)

Kp= P[NH3]XP[HI]= 137.5 X 137.5 ---------(2)
For 2HI H2+ I2
1 00
(1 - x)x/2 x/2

2HI H2+ I2 Kp= (0.015)2x = 0.2

Again since HI decomposes reaction proceeds in forward directioal.

But a = (P + A) X 0.2
also kp
for final eqn (3)
= 0.8(P + A)2But Kp
= 137.5 X 137.5
0.8(P + A)2
= 137.5 X 137.5
P + A = 153.73
A = 16.22

a = (p + A) X 0.2 = 153.73 X 0.2 = 30.75
Total pressure at eqn =P[NH3]+P[HI]+P[H2]+P[I]=(P+A)+(P+A-a)+A/2+A/2=2p+2A
=2(p+A)= 2x153.73
=307.46 mmHg
Akshita Goyal
8 years ago
can u pls specify  what is A and a and how did the equatn nh4i forms nh3 (P+A)+hI(P+A-a) come?